2) -n-9+8n=26
3) 5(w-7)=-15
4) m=3(m-4)=16 I NEED THIS QUESTIONS PLEASE

Answer by stanbon(75887)

2) -n-9+8n=26
7n-9=26
7n=35
n=5
3) 5(w-7)=-15
5w-35=-15
5x=20
x=4
4) m=3(m-4)+16
m=3m-12+16
-2x=4
x=-2
Cheers,
Stan H.